Blitz Industries has a debt-equity ratio of 1.6. Its WACC is 7.8 percent, and its cost of debt is 5.5 percent. The corporate tax rate is 21 percent. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

a. What is the company’s cost of equity capital?

b. What is the company’s unlevered cost of equity capital?

c-1. What would the cost of equity be if the debt-equity ratio were 2?

c-2. What would the cost of equity be if the debt-equity ratio were 1.0?

c-3. What would the cost of equity be if the debt-equity ratio were zero?
Answer #1

a)

WACC = 7.8% Cost of debt (rd) = 5.5% Cost of equity (re) = ? Tax rate (T) =21% Debt equity ratio = 1.6 WACC = Wd* rd *(1-T) +

b)

re=r0+ (ro-rd) * (1-t) * (d/e) re = required return on equity ro = required return on equity with zero debt t = tax rate d/e

c-1)

re=r0 + (ro-rd) *(1-t) * (d/e) re = required return on equity ro = required return on equity with zero debt t = tax rate d/e

c-2)

  re=r0+ (ro-rd) *(1-t) * (d/e) re = required return on equity ro = required return on equity with zero debt t = tax rate d/e =

c-3)

  re=r0 + (ro-rd) * (1-t) * (d/e) re = required return on equity ro = required return on equity with zero debt t = tax rate d/e